# Towards Operational Automated Greenhouse Gas Plume Detection

### v1

@@ Abstract
Operational deployment of a fully automated greenhouse gas (GHG) plume detection system remains an elusive goal for imaging spectroscopy missions, despite recent advances in deep learning approaches. With the dramatic increase in data availability, however, automation continues to increase in importance for natural and anthropogenic emissions monitoring. This work reviews and addresses several key obstacles in the field: data and label quality control, prevention of spatiotemporal biases, and correctly aligned modeling objectives. We demonstrate through rigorous experiments using multicampaign data from airborne and spaceborne instruments that convolutional neural networks (CNNs) are able to achieve operational detection performance when these obstacles are alleviated. We demonstrate that a multitask model that learns both instance detection and pixelwise segmentation simultaneously can successfully lead towards an operational pathway. We evaluate the model’s plume detectability across emission source types and regions, identifying thresholds for operational deployment. Finally, we provide analysis-ready data, models, and source code for reproducibility, and work to define a set of best practices and validation standards to facilitate future contributions to the field.
@@
Plume Detection , Greenhouse Gas , Machine Learning , Imaging Spectroscopy[][nocite]build/opsghgml_supplemental
[jpl]organization=Jet Propulsion Laboratory, California Institute of Technology, city=Pasadena, state=CA, postcode=91101, country=USA
@@ 1 Introduction
Modern airborne & spaceborne imaging spectrometers such as the Airborne Visible-Infrared Imaging Spectrometer-Next Generation (AVIRIS-NG) (Hamlin et al., 2011; Chapman et al., 2019), the Global Airborne Observatory (GAO) (Asner, 2007; Asner et al., 2012), the Earth surface Mineral dust source InvesTigation (EMIT) (Green et al., 2020), and Carbon Mapper Coalition’s Tanager-1 (Zandbergen et al., 2023) capture finely-sampled observations with spectral resolution sufficient to detect GreenHouse Gas (GHG) emissions with discriminative absorptions in the Visible to ShortWave InfraRed (VSWIR) and spatial resolution sufficient to resolve meter-scale surface infrastructure elements.
@@
Spatially-informed supervised machine learning models such as Convolutional Neural Networks (CNNs) provide a potential solution to automatically detect, localize, and delineate plumes from GHG retrievals. CNN-based approaches have proven successful in a variety of complex image classification, object detection, and pixelwise segmentation benchmark data sets (Russakovsky et al., 2014; Everingham et al., 2010; Krizhevsky, 2009). Recent studies have claimed impressive results using CNN-based plume detection systems driven by GHG image products derived from data captured by airborne and spaceborne instruments (Joyce et al., 2022; Schuit et al., 2023; Bruno et al., 2023; Vaughan et al., 2023; Kumar et al., 2023; Rouet-Leduc and Hulbert, 2024).
@@
We address these issues in this work and in our previous efforts Bue et al. (2023, 2024); Lee et al. (2021, 2022, 2023) by guided by the following research questions:
@@
This paper is organized as follows: Section 2 provides background information on the CMF retrieval product, GHG plumes and their labels, and definitions of the plume detection task. Section 3 describes the plume detection system, including quality control, sampling, and validation. Section 4 describes this work’s data sets and models, and Section 5 presents the corresponding plume detection results. Finally, Section 7 puts these results in context of limitations in existing literature, and we conclude in Section 8.
@@ 2 Background
To define operational requirements for a plume detection system, we first must understand the data and labels that will be used to construct and validate the system. Equipped with this knowledge, we can effectively construct protocols to collect representative samples to train and test the model, define optimization routines that account for task-specific sources of bias and uncertainty, and select capable validation metrics that provide realistic estimates of generalization performance. While the data and results described in this work are focused on matched filter-based retrievals of methane derived from imaging spectrometer observations, the challenges we describe are generally relevant to similar plume detection systems, including those based on image data captured by analogous remote sensing platforms such as PRISMA and EnMAP (Guanter et al., 2023; Roger et al., 2024), targeting other greenhouse gases such as or , and retrievals derived using alternative algorithms such as mag1c (Foote et al., 2020).
@@ 2.1 GHG Retrieval Image Products
We derive pixelwise GHG concentrations from calibrated imaging spectrometer radiance observations using the Columnwise Matched Filter (CMF) (Thompson et al., 2015) retrieval approach. Based largely on the methods developed and refined by scientists at Los Alamos National Labs starting in the late 1990s (Villeneuve et al., 1999; Theiler, 2012), the CMF is conceptually simple, computationally efficient, and produces retrievals that are both numerically stable across repeat observations of the same sites and emission sources and consistent with coincident measurements by complementary sensors (e.g., comparisons with dynamic aviation soundings collected during Four Corners campaign (Thompson et al., 2015)) and with measurements from controlled release experiments (Thorpe et al., 2017, 2023, 2016; Duren et al., 2019; El Abbadi et al., 2024; Sherwin et al., 2023; Ayasse et al., 2023). The CMF retrieval computes the pixelwise enhancement of VSWIR absorption features associated with a MODTRAN-simulated GHG transmittance spectrum in units relative to multivariate Gaussian backgrounds, each estimated columnwise from the observations of the independent detector elements in the sensor’s focal plane array. CMF products derived from AVIRIS-NG and GAO data are highly analogous—both are based on the approach described in Thompson et al. (2015) that retrieves pixelwise concentrations according to the depth of the SWIR2 absorption feature of airborne spectrometer radiance observations captured at GSD [1,10] m. Any differences between the instruments are dwarfed by atmospheric, geometric, and radiative transfer assumptions within the retrieval. EMIT observations are observed at 60 m GSD and EMIT CMF products are derived using the same matched-filter-based algorithm used with the airborne data. The operational EMIT L2 product uses all VSWIR wavelengths rather than just the SWIR2, is more stable due to its environment in orbit, and leverages a Look-Up Table (LUT) to select an appropriate target spectrum for the retrieval that accounts for varying concentration levels and solar zenith angles (Foote et al., 2021; Thorpe et al., 2023). Filtering of cloud, water, and thermal anomalies from the calculation of the covariance also avoids artifacts in the retrieval.
@@ 2.3 Plume Labels
This work leverages plumes identified by domain experts as described in Suppl. Sections LABEL:apx:qc and LABEL:apx:airborne. Constructing and validating a ML-driven GHG plume detector also requires pixelwise labels to teach the model to distinguish plume pixels from background pixels. However, plumes are not discrete objects with well-defined boundaries, and manually labeling individual plume pixels is a difficult and subjective task, particularly near plume boundaries. One alternative to requiring users to manually label individual plume pixels is to use the CMF enhancements to deterministically generate “CMF-guided” plume labels consisting of highly concentrated pixels near the origin point of each user identified plume. While the CMF-guided labels may be less subjective than manually defined labels, their quality can vary significantly by scene according to the local contrast in values between plumes and background enhancements or the presence of false enhancements adjacent to real plumes, and consequently require comprehensive manual inspection to refine or reject uninformative instances.
Figure 2 shows two example plume candidates (left column) with corresponding labels provided by four independent experts (center column), along with the deterministic CMF-guided plume labels for each plume according to increasing thresholds (right column). All four users agree that each enhancement represents a real plume, but the pixelwise masks they each provide disagree with respect to the boundaries where plume pixels end and background pixels begin (center column). In contrast, the CMF-guided labels are more conservative, capturing only the highly concentrated pixels near the origin point associated with each plume.
@@ 2.4 Background and False Enhancements
Background Enhancements: The majority of pixels in a typical scene are common background enhancements representing retrievals of surface materials lacking characteristic GHG absorption features varying with scenewise imaging geometry, albedo, illumination conditions, and atmospheric state. Most background enhancements exhibit small values that fall within the range where GHG concentrations are too low to distinguish them from instrument noise (e.g., 250 in typical airborne CMFs). By definition, the CMF column concentration should follow a zero-mean distribution, with the diffuse background emerging as a consequence of instrument noise characteristics and environmental conditions. Though higher concentration retrievals occasionally occur in regions dominated by common background enhancements, they tend to be spatially sparse. Consequently, regions consisting primarily of common background enhancements consist of mostly low concentration pixels with sporadic high enhancement pixels that lack cohesive spatial structure, making such regions easily distinguishable from real GHG plumes.
False Enhancements: False enhancements are a nuisance class of background enhancements with concentration levels within the range of real GHG plumes (exceeding 500-1000) caused by factors unrelated to the presence of GHG emissions. One common driver involves confusers, including surface spectroscopy that mimics methane absorption features at the spectral sampling interval of the instrument and anomalous observations, such as high albedo materials and thermal anomalies. Another involves artifacts caused by limitations of the instrument or retrieval method, such as corrupt background covariance estimates due to flares and sample size limitations from short flightlines. Although false enhancements are relatively rare in terms of pixelwise coverage (we estimate false enhancements represent 1-5% of the retrieved pixels in a typical CMF scene) false enhancements that manifest at spatial scales and with spatial structure similar to real plumes constitute the majority of false positive plume detections observed in this work, and can be difficult even for domain experts to distinguish from real GHG plumes. False enhancements can severely restrict plume detectability when they occur near real emission sources and merge with pixels representing observed plumes, thereby altering their spatial distribution and concentration levels. Additionally, in some rare, pathological cases, false enhancements can cover the spatial extent of an entire scene (see examples in Suppl. Section LABEL:apx:gao_penn and Suppl. Figure LABEL:fig:penn_tiles).
@@ 2.5 Operational Plume Detection
Both of these objectives are achievable using current GHG image retrieval approaches applied to data captured by modern airborne and spaceborne imaging spectrometer platforms. Prior manual plume identification efforts have shown that human experts can consistently identify, localize and delineate most plumes through visual inspection of the CMF retrieval images we consider in this work. Comprehensive manual analysis of retrieval images captured in airborne imaging campaigns has also informed improvements to GHG retrieval algorithms, suppressing common classes of false enhancements, and providing diagnostic capabilities to quantify retrieval sensitivity and uncertainty.
@@ 3 Methods
System Overview: We consider plume detection systems of the form shown in Figure 4 in this work. Provided a set of labeled scenes consisting of GHG retrieval image products and their associated pixelwise plume labels, our objective is to train a plume detector model to estimate the probability (or “plume salience”) that each pixel in the image represents a plume. Supervised machine learning models achieve this by learning to map labeled plume image pixels to high salience values and unlabeled “background” image pixels to low salience values. In this work, we focus on CNN-based plume detectors which learn a hierarchy of spatial filters that capture disscriminative features of labeled plumes and unlabeled background enhancements using fixed-size image tiles extracted from our labeled CMF scenes. Once trained, CNNs are capable of efficiently generating spatially-informed pixelwise predictions (or “plume salience maps”) based on CMF images with arbitrary input dimensions.
@@
To ensure our models and results are not spatiotemporally biased, we use the bounding boxes of the georeferenced CMF images to stratify our data into spatially disjoint training and test scenes. For airborne campaigns, we compute the graph of pairwise Intersection-over-Union (IoU) scores with respect to the scenes observed in each campaign. Each connected component of the resulting IoU graph identifies a group of overlapping flightlines that are spatially disjoint from all the remaining flightlines. For the spaceborne EMIT data, we group scenes according to their respective UTM & MGRS zone ids. We generate spatially disjoint, class stratified training and test partitions by assigning the labeled plume and background tiles associated with each group of flightlines to either the training set or the test set via the GroupShuffleSplit function provided in the scikit-learn package (Pedregosa et al., 2011). For each campaign, we designate roughly 75% of the flightlines (and their associated tiles) as training data, and the remaining 25% of the flightlines as test data.
@@ 4.2 Airborne Plume Detector Data
Unique Campaigns: CMF image products derived from data captured in the 2015 AVIRIS-NG Four Corners campaign and the 2021 GAO Penn campaign exhibit characteristics that are distinct from CMF images derived from data captured in other airborne campaigns. We summarize how these distinctions impact plume detection capabilities and the interpretation of plume detection result in Suppl. Section LABEL:apx:bias_region, but we exclude the observations from these anomalous campaigns from all of our remaining experiments.
@@ 4.3 EMIT Plume Detector Data
EMIT VISIONS Data Set: All plume detection experiments involving EMIT observations were trained and tested on observations from the EMIT VISIONS Data Set. The EMIT VISIONS data set consists of the expert curated, high SnR CMF image products derived from data captured by the spaceborne imaging spectrometer EMIT between 2022-2024 that contain high-confidence methane plumes delivered to the public-facing EMIT VISIONS portal 111https://earth.jpl.nasa.gov/emit/data/data-portal/Greenhouse-Gases/ that were formally-reviewed & approved by multiple reviewers using the Multi-Mission Geographic Information System (Soliman et al., 2025).

### v2

@@ Abstract
Operational deployment of a fully automated facility-scale greenhouse gas (GHG) plume detection system remains challenging for fine spatial resolution imaging spectrometers, despite recent advances in deep learning approaches. With the dramatic increase in data availability, however, automation continues to increase in importance for emissions monitoring. This work reviews and addresses several key obstacles in the field: data and label quality control, prevention of spatiotemporal biases, and correctly aligned modeling objectives. We demonstrate through rigorous experiments using multicampaign data from airborne and spaceborne instruments that convolutional neural networks (CNNs) are able to achieve operational detection performance when these obstacles are alleviated. We demonstrate that a multitask model that learns both instance detection and pixelwise segmentation simultaneously can successfully lead towards an operational pathway. We evaluate the model’s plume detectability across emission source types and regions, identifying thresholds for operational deployment. Finally, we provide analysis-ready data, models, and source code for reproducibility, and work to define a set of best practices and validation standards to facilitate future contributions to the field.
@@
Plume Detection , Greenhouse Gas , Machine Learning , Imaging Spectroscopy , AVIRIS , EMIT[jpl]organization=Jet Propulsion Laboratory, California Institute of Technology, city=Pasadena, state=CA, postcode=91101, country=USA
@@ 1 Introduction
Modern airborne & spaceborne imaging spectrometers such as the Airborne Visible-Infrared Imaging Spectrometer-Next Generation (AVIRIS-NG) (Hamlin et al., 2011; Chapman et al., 2019), the Global Airborne Observatory (GAO) (Asner, 2007; Asner et al., 2012), the Earth surface Mineral dust source InvesTigation (EMIT) (Green et al., 2020), and Carbon Mapper Coalition’s Tanager-1 (Zandbergen et al., 2023) capture finely-sampled observations with spectral resolution sufficient to detect GreenHouse Gas (GHG) emissions with discriminative absorptions in the Visible to ShortWave InfraRed (VSWIR) and spatial resolution sufficient to attribute emissions to specific infrastructure elements.
@@
Spatially-informed supervised machine learning models such as Convolutional Neural Networks (CNNs) provide a potential solution to automatically detect, localize, and delineate plumes from GHG retrievals. CNN-based approaches have proven successful in a variety of complex image classification, object detection, and pixelwise segmentation benchmark data sets (Russakovsky et al., 2015; Everingham et al., 2010; Krizhevsky, 2009). Recent studies have claimed impressive results using CNN-based plume detection systems driven by GHG image products derived from data captured by airborne and spaceborne instruments (Joyce et al., 2022; Schuit et al., 2023; Bruno et al., 2024; Vaughan et al., 2024; Kumar et al., 2023; Rouet-Leduc and Hulbert, 2024).
@@
We address these issues in this work and in our previous efforts (Bue et al., 2023, 2024; Lee et al., 2021, 2022, 2023) guided by the following research questions:
@@
This paper is organized as follows: Section 2 provides background information on the Columnwise Matched Filter (CMF) retrieval product, GHG plumes and their labels, and definitions of the plume detection task. Section 3 describes the plume detection system, including quality control, sampling, and validation. Section 4 describes this work’s data sets and models, and Section 5 presents the corresponding plume detection results. Finally, Section 6 puts these results in context of limitations in existing literature, and we conclude in Section 7.
@@ 2 Background
To define operational requirements for a plume detection system, we first must understand the data and labels that will be used to construct and validate the system. Equipped with this knowledge, we can effectively construct protocols to collect representative samples to train and test the model, define optimization routines that account for task-specific sources of bias and uncertainty, and select capable validation metrics that provide realistic estimates of generalization performance. While the data and results described in this work are focused on matched filter-based retrievals of methane derived from imaging spectrometer observations, the challenges we describe are generally relevant to similar plume detection systems, including those based on image data captured by analogous remote sensing platforms such as PRISMA and EnMAP (Guanter et al., 2023; Roger et al., 2024), targeting other greenhouse gases such as , (Borger et al., 2025) or (Balasus et al., 2026), and retrievals derived using alternative algorithms such as mag1c (Foote et al., 2020).
@@ 2.1 GHG Retrieval Image Products
We derive pixelwise GHG concentrations from calibrated imaging spectrometer radiance observations using the Columnwise Matched Filter (CMF) (Thompson et al., 2015) retrieval approach. Based largely on the methods developed and refined by scientists at Los Alamos National Labs starting in the late 1990s (Villeneuve et al., 1999; Theiler, 2012), the CMF is conceptually simple, computationally efficient, and produces retrievals that are both numerically stable across repeat observations of the same sites and emission sources and consistent with coincident measurements by complementary sensors (e.g., comparisons with Dynamic Aviation in situ observations collected during Four Corners campaign (Thompson et al., 2015)) and with measurements from controlled release experiments (Thorpe et al., 2017, 2023, 2016; Duren et al., 2019; El Abbadi et al., 2024; Sherwin et al., 2023; Ayasse et al., 2023). The CMF retrieval computes the pixelwise enhancement of VSWIR absorption features associated with a MODTRAN-simulated GHG transmittance spectrum in ppm-m units relative to multivariate Gaussian backgrounds, each estimated columnwise from the observations of the independent detector elements in the sensor’s focal plane array. CMF products derived from AVIRIS-NG and GAO data are highly analogous—both are based on the approach described in Thompson et al. (2015) that retrieves pixelwise concentrations according to the depth of the SWIR2 absorption feature of airborne spectrometer radiance observations captured at GSD [1,10] m. Any differences between the instruments are dwarfed by atmospheric, geometric, and radiative transfer assumptions within the retrieval. EMIT observations are observed at 60 m GSD and EMIT CMF products are derived using the same matched-filter-based algorithm used with the airborne data. The operational EMIT L2 product uses all VSWIR wavelengths rather than just the SWIR2, is more stable due to its environment in orbit, and leverages a Look-Up Table (LUT) to select an appropriate target spectrum for the retrieval that accounts for varying concentration levels and solar zenith angles (Foote et al., 2021; Thorpe et al., 2023). Filtering of cloud, water, and thermal anomalies from the calculation of the covariance also avoids artifacts in the retrieval.
@@ 2.3 Plume Labels
This work leverages plumes identified by domain experts as described in Suppl. Sections LABEL:apx:qc and LABEL:apx:airborne. Constructing and validating a ML-driven GHG plume detector also requires pixelwise labels (binary masks) to teach the model to distinguish plume pixels from background pixels. However, plumes are not discrete objects with well-defined boundaries, and manually labeling individual plume pixels is a difficult and subjective task, particularly near plume boundaries. One alternative to requiring users to manually label individual plume pixels is to use the CMF enhancements to deterministically generate “CMF-guided” plume labels consisting of highly concentrated pixels near the origin point of each user identified plume. While the CMF-guided labels may be less subjective than manually defined labels, their quality can vary significantly by scene according to the local contrast in ppm-m values between plumes and background enhancements or the presence of false enhancements adjacent to real plumes, and consequently require comprehensive manual inspection to refine or reject uninformative instances.
Figure 2 shows two example plume candidates (left column) with corresponding labels provided by four independent experts (center column), along with the deterministic CMF-guided plume labels for each plume according to increasing ppm-m thresholds (right column). The experts’ labels are smoother and coarser compared to the CMF-guided plume labels, as they are defining polygons by clicking on vertices. Experts were also instructed to include a small buffer region beyond the diffused edges of the plume to capture the entire extent. All four users agree that each enhancement represents a real plume, but the pixelwise masks they each provide disagree with respect to the boundaries where plume pixels end and background pixels begin (center column). In contrast, the CMF-guided labels are more conservative, capturing only the highly concentrated pixels near the origin point associated with each plume. The CMF-guided labeling process is described in detail in Suppl. Section LABEL:sub:cmflab.
@@ 2.4 Background and False Enhancements
Background Enhancements: The majority of pixels in a typical scene are common background enhancements representing retrievals of surface materials lacking characteristic GHG absorption features varying with scenewise imaging geometry, albedo, illumination conditions, and atmospheric state. Most background enhancements exhibit small values that fall within the range where GHG concentrations are too low to distinguish them from instrument noise (e.g., 250 ppm-m in typical airborne CMFs). By definition, the CMF column concentration should follow a zero-mean distribution, with the diffuse background emerging as a consequence of instrument noise characteristics and environmental conditions. Though higher concentration retrievals occasionally occur in regions dominated by common background enhancements, they tend to be spatially sparse. Consequently, regions consisting primarily of common background enhancements consist of mostly low concentration pixels with sporadic high enhancement pixels that lack cohesive spatial structure, making such regions easily distinguishable from real GHG plumes.
False Enhancements: False enhancements are a nuisance class of background enhancements with concentration levels within the range of real GHG plumes (exceeding 500-1000ppm-m) caused by factors unrelated to the presence of GHG emissions. One common driver involves confusers, including surface spectroscopy that mimics methane absorption features at the spectral sampling interval of the instrument and anomalous observations, such as high albedo materials and thermal anomalies. Another involves artifacts caused by limitations of the instrument or retrieval method, such as corrupt background covariance estimates due to flares and sample size limitations from short flightlines. Although false enhancements are relatively rare in terms of pixelwise coverage (we estimate false enhancements represent 1-5% of the retrieved pixels in a typical CMF scene) false enhancements that manifest at spatial scales and with spatial structure similar to real plumes constitute the majority of false positive plume detections observed in this work, and can be difficult even for domain experts to distinguish from real GHG plumes. False enhancements can severely restrict plume detectability when they occur near real emission sources and merge with pixels representing observed plumes, thereby altering their spatial distribution and concentration levels. Additionally, in some rare, pathological cases, false enhancements can cover the spatial extent of an entire scene (see examples in Suppl. Section LABEL:apx:gao_penn and Suppl. Figure LABEL:fig:penn_tiles).
@@ 2.5 Operational Plume Detection
Both of these objectives are achievable using current GHG image retrieval approaches applied to data captured by modern airborne and spaceborne imaging spectrometer platforms. Prior manual plume identification efforts have shown that human experts can consistently identify, localize, and delineate most plumes above the minimum detection limit of EMIT through visual inspection of the CMF retrieval images we consider in this work (Duren et al., 2019; Thorpe et al., 2023).
Comprehensive manual analysis of retrieval images captured in airborne imaging campaigns has also informed improvements to GHG retrieval algorithms, suppressing common classes of false enhancements, and providing diagnostic capabilities to quantify retrieval sensitivity and uncertainty.
@@ 3 Methods
System Overview: We consider plume detection systems of the form shown in Figure 4 in this work. Provided a set of labeled scenes consisting of GHG retrieval image products and their associated pixelwise plume labels, our objective is to train a plume detector model to estimate the probability (or “plume salience”) that each pixel in the image represents a plume. Supervised machine learning models achieve this by learning to map labeled plume image pixels to high salience values and unlabeled “background” image pixels to low salience values. In this work, we focus on CNN-based plume detectors which learn a hierarchy of spatial filters that capture discriminative morphological features of labeled plumes and unlabeled background enhancements using fixed-size image tiles extracted from our labeled CMF scenes. Once trained, CNNs are capable of efficiently generating spatially-informed pixelwise predictions (or “plume salience maps”) based on CMF images with arbitrary input dimensions.
@@
To ensure our models and results are not spatiotemporally biased, we use the bounding boxes of the georeferenced CMF images to stratify our data into spatially disjoint training and test scenes. For airborne campaigns, we compute the graph of pairwise Intersection-over-Union (IoU) scores with respect to the scenes observed in each campaign. Each connected component of the resulting IoU graph identifies a group of overlapping flightlines that are spatially disjoint from all the remaining flightlines. For the spaceborne EMIT data, we group scenes according to their respective Universal Transverse Mercator (UTM) zone IDs. We generate spatially disjoint, class stratified training and test partitions by assigning the labeled plume and background tiles associated with each group of flightlines to either the training set or the test set via the GroupShuffleSplit function provided in the scikit-learn package (Pedregosa et al., 2011). For each campaign, we designate roughly 75% of the flightlines (and their associated tiles) as training data, and the remaining 25% of the flightlines as test data.
@@ 4.2 Airborne Plume Detector Data
Unique Campaigns: CMF image products derived from data captured in the 2015 AVIRIS-NG Four Corners campaign and the 2021 GAO Pennsylvania (Penn) campaign exhibit characteristics that are distinct from CMF images derived from data captured in other airborne campaigns. We summarize how these distinctions impact plume detection capabilities and the interpretation of plume detection result in Suppl. Section LABEL:apx:bias_region, but we exclude the observations from these anomalous campaigns from all of our remaining experiments.
@@ 4.3 EMIT Plume Detector Data
EMIT VISIONS Data Set: All plume detection experiments involving EMIT observations were trained and tested on observations from the EMIT VISIONS Data Set. The EMIT VISIONS data set consists of the expert curated, high SnR CMF image products derived from data captured by the spaceborne imaging spectrometer EMIT between 2022-2024 that contain high-confidence methane plumes delivered to the public-facing EMIT VISIONS portal 111https://earth.jpl.nasa.gov/emit/data/data-portal/Greenhouse-Gases/ that were formally-reviewed & approved by multiple reviewers using the Multi-Mission Geographic Information System (Soliman et al., 2025). For experiments with EMIT data, we use the readily available expert plume extent delineation products as segmentation mask labels instead of CMF-guided labels.